# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ B.2.2 Finite-sum setting
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$. We want the right hand side of both bounds $\to 0$. In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
$\min_{p,w:0<p\leq\omega}\max\left\{t^{p-1},t^{w-1},t^{p-w},t^{-w},t^{-p},t^{1-2w},t^{1-w-p}\right\}=-\frac{1}{4},$
which is realized by setting $p=1/2$, $\omega=3/4$ as required. ∎
B.3 Additional numerical experiments
We present an additional numerical for a dictionary learning problem to illustrate the performance of Algorithms 1 and 2 in the nonconvex setting. We highlight that, similar to Theorems 3 and 4, one can show from Theorems 5 and 6 that there exists a subsequence of the iterates $\{x_{t}\}_{t\geq 0}$ generated by Algorithms 1 and 2 that converges almost surely to a stationary point of Eq. 1. This follows from the fact that the auxiliary function $\mathcal{F}$ can be shown to be Lipschitz continuous. Moreover, using the same argument as in the proof of [4, Theorem 4.5 and 4.7], it is possible to show that there exists a point $x_{t^{\star}}$ in a subsequence of $\{x_{t}\}_{t\geq 0}$ that satisfies a stationarity guarantee similar to that in Theorems 5 and 6. Since this is merely an existence result and the proof is long, we omit the details for brevity. Nonetheless, we present the convergence behavior of the last iterate of the sequence $\{x_{t}\}_{t\geq 0}$ in a dictionary learning problem, which is also studied in [4, 28].
$\min_{D,X}&\quad\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Dx_{i}\|_{2}^{2}\\ \text{s.t.}~{}&\quad D\in{\mathbb{R}}^{m\times p},\ X=[x_{1}\cdots x_{n}]\in{\mathbb{R}}^{p\times n}\\ &\quad\|d_{j}\|_{2}\leq 1,\quad\forall j\in[p]\\ &\quad\|x_{i}\|_{1}\leq\delta,\quad\forall i\in[n],\end{aligned}$
where we refer to $X$ as the coefficient matrix.
In practice, data points usually arrive sequentially and the representation evolves gradually. Hence, the dictionary must be updated sequentially as well. Assume that we already have learned a dictionary $\hat{D}\in{\mathbf{R}}^{m\times p}$ and the corresponding coefficient matrix $\hat{X}\in{\mathbf{R}}^{p\times n}$ for some data set $A$. As a new dataset $A^{\prime}=\{a_{1}^{\prime},\dots,a_{n^{\prime}}^{\prime}\}\subseteq{\mathbf{R}}^{m}$ arrives, we intend to enrich our dictionary by learning $\tilde{D}\in{\mathbf{R}}^{m\times q}$ ($q>p$) and the coefficient matrix $\tilde{X}\in{\mathbf{R}}^{q\times n^{\prime}}$ for the new dataset while maintaining good performance of $\tilde{D}$ on the old dataset $A$ as well as the learned coefficient matrix $\hat{X}$ This leads to the following finite-sum bilevel problem:
$\min_{\tilde{D},\tilde{X}}$ $\quad F(\tilde{D},\tilde{X}):=\frac{1}{2n^{\prime}}\sum_{i\in[n^{\prime}]}\|a_{i}^{\prime}-\tilde{D}\tilde{x}_{i}\|_{2}^{2}$ (32)
s.t. $\quad\tilde{D}\in{\mathbb{R}}^{m\times q},\ \tilde{X}=[\tilde{x}_{1}\cdots\tilde{x}_{n}]\in{\mathbb{R}}^{q\times n^{\prime}}$
$\quad\|\tilde{x}_{i}\|_{1}\leq\delta,\quad\forall i\in[n^{\prime}],$
$\quad\tilde{D}\in\operatorname*{arg\,min}_{\|z_{j}\|_{2}\leq 1,\forall j\in[q]}g(Z):=\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Z\hat{x}_{i}\|_{2}^{2},$
where we denote $\hat{x}_{i}$ as the prolonged vector in ${\mathbf{R}}^{q}$ by appending zeros at the end.
In our experiment, we obtain $A,A^{\prime},\hat{X}$ from the code provided by Cao et al. [4] with $n=n^{\prime}=250$ and $p=40,q=50$ In addition, we also choose $\delta=3$. The experiment is repeated $10$ times, and we report the average performance within a $1$-minute execution time limit. For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-2/7}$ and $\alpha_{t}=(t+1)^{-6/7}$, where $\varsigma=0.1$. For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-1/2}$, where $\varsigma=0.1$, along with $\alpha_{t}=(t+1)^{-3/4}$ for every $t\geq q+1$ and $\alpha_{t}=\log(q+1)/(q+1)$ for every $t\leq q$. In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.1(t+1)^{-2/3}$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-3}$ (both using $K_{t}=0.01(t+1)^{-1/3},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [26] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-4}/(t+1)^{3/4}$, $\rho_{t}=(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [20] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=5\times 10^{-3}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{2}$-norm ball intersecting with a half-space is required, which we employ CVXPY [10] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1] to compute projection on $\ell_{1}$ norm ball and projection onto $\ell_{2}$ norm ball admits an analytical solution. To approximate inner-level optimal value $G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem.
Figure 3 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) problems in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves better solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4]. We also observe that SDBGD and IR-SCG perform less number of iterations compared to our methods.
Method Average # of iterations Average # of oracle calls
IR-SCG $32736.1$ $65472.2$
IR-FSCG $30532.6$ $1872851.0$
SBCGI $466.3$ $7440.6$
SBCGF $361.84$ $6147916.0$
aR-IP-SeG (short step) $6290.0$ $25160.0$
aR-IP-SeG (long step) $6278.1$ $25112.4$
SDBGD (short step) $11441.3$ $22882.6$
SDBGD (long step) $12239.3$ $24478.6$
B.4 Finiteness of constants
@@
If 1 and 1 to 3 hold, then $C,\bar{C}_{\delta},C_{\delta}$ defined in (15) are finite. Furthermore, if $L$ defined in 2 is strictly less than $1$, then $V$ defined in (15d) is finite.
Proof of Lemma 16.
We first show that under 1 and 3, when the limit in 2 exists, the parameter $L$ satisfies $0\leq L\leq 1$. While it is trivial to see $L\geq 0$ thanks to 1, $L\leq 1$ needs some justifications. For the sake of contradiction, suppose $L>1$. Then for some sufficiently large $t$, we have
@@
which always contradicts 3. Thus, we have $0\leq L\leq 1$ under our blanket assumptions.
@@
where the first equality is obtained by diving both the numerator and the denominator by $t\sigma_{t}$ and the second equality follows from 2. Moreover, we have
@@
Using 2, we have $\min_{t\geq 0}(t+1)\sigma_{t}>0$. Thus, $C$ is finite. It is also straightforward to verify that $\bar{C}_{\delta}$ is finite by 3. Consequently, $C_{\delta}$ is finite as well.
To conclude, we show $V$ is finite. From 3, $\{t\sigma_{t}\}_{t\geq 0}$ and $\{(t+1)t\sigma_{t}^{2}\}_{t\geq 0}$ are increasing and diverge to $\infty$. Note that
@@
where the first equality is implied by the Stolz-Cesàro theorem, the second equality is obtained from dividing both the denominator and the numerator by $(t+1)\sigma_{t+1}^{2}$, and the third equality comes from 3 and
@@
If 1, 1, 2 and 4 hold, then $C_{q},\bar{C}_{\delta,q}$ defined in (23f), and (23g) are finite. Furthermore, if $L$ defined in 2 is less than $1$, then $V_{q}$ defined in (23h) is finite.
Proof of Lemma 17.
The proof is similar to the proof of Lemma 16. Details are omitted for brevity. ∎

### v2

@@ C.2 Proof of Theorem˜4
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability at least $1-\delta$. We want the right hand side of both bounds $\to 0$. In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
$\min_{p,w:0<p\leq\omega}\max\left\{p-1,w-1,p-w,-w,-p,1-2w,1-w-p\right\}=-\frac{1}{4},$
which is realized by setting $p=1/2$, $\omega=3/4$ as required.
As for the second claim on asymptotic convergence, we prove a more general result: If
$\max\left\{p-1,w-1,p-w,-w,-p,1-2w,1-w-p\right\}<0,$
then the asymptotic convergence holds. Under this additional assumption and the fact that $\mathcal{G}(x_{t})\geq 0$, we deduce that $\lim_{t\to\infty}\mathcal{G}(x_{t})=0$ from taking the limit on both sides of (32). Recall that
$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\inf_{k\geq t}\mathcal{F}(x_{k}),$
and
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\geq\frac{\sum_{i=0}^{t}\beta_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\beta_{i}}.$
Combining these observations and (33), we have
$\frac{\sum_{i=0}^{t}\beta_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\beta_{i}}$
$\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}\left(a+bt^{1-p}\right)^{-1}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1}.$
By the Stolz–Cesàro theorem (since $p\in(0,1)$, $\sum_{t\geq 0}\beta_{t}=\infty$) and taking the limit on both sides of the above inequality, we obtain $\liminf_{t\to\infty}\mathcal{F}(x_{t})\leq 0$. Since this holds with probability at least $1-\delta$ for any $\delta\in(0,1)$, it also holds almost surely. Due to the fact that any limit point of $\{x_{t}\}_{t}$ is in $X_{\operatorname{opt}}$, the continuity of $\mathcal{F}$ from Lemma˜4 and the fact that $\mathcal{F}(x)\geq 0$ for any $x\in X_{\operatorname{opt}}$, we deduce that $\liminf_{t\to\infty}\mathcal{F}(x_{t})\geq 0$. Thus, we conclude the proof. ∎
Appendix D Additional Discussions
D.1 Finiteness of constants
@@
If ˜1 and 1 to 3 hold, then $C,\bar{C}_{\delta},C_{\delta}$ defined in (15) are finite. Furthermore, if $L$ defined in ˜2 is strictly less than 11, then $V$ defined in (15d) is finite.
Proof of Lemma˜17.
We first show that under ˜1 and ˜3, when the limit in ˜2 exists, the parameter $L$ satisfies $0\leq L\leq 1$. While it is trivial to see $L\geq 0$ thanks to ˜1, $L\leq 1$ needs some justifications. For the sake of contradiction, suppose $L>1$. Then for some sufficiently large $t$, we have
@@
which always contradicts ˜3. Thus, we have $0\leq L\leq 1$ under our blanket assumptions.
@@
where the first equality is obtained by diving both the numerator and the denominator by $t\sigma_{t}$ and the second equality follows from ˜2. Moreover, we have
@@
Using ˜2, we have $\min_{t\geq 0}(t+1)\sigma_{t}>0$. Thus, $C$ is finite. It is also straightforward to verify that $\bar{C}_{\delta}$ is finite by ˜3. Consequently, $C_{\delta}$ is finite as well.
To conclude, we show $V$ is finite. From ˜3, $\{t\sigma_{t}\}_{t\geq 0}$ and $\{(t+1)t\sigma_{t}^{2}\}_{t\geq 0}$ are increasing and diverge to $\infty$. Note that
@@
where the first equality is implied by the Stolz-Cesàro theorem, the second equality is obtained from dividing both the denominator and the numerator by $(t+1)\sigma_{t+1}^{2}$, and the third equality comes from ˜3 and
@@
If ˜1, ˜1, 2 and 4 hold, then $C_{q},\bar{C}_{\delta,q}$ defined in (26f), and (26g) are finite. Furthermore, if $L$ defined in ˜2 is less than 11, then $V_{q}$ defined in (26h) is finite.
Proof of Lemma˜18.
The proof follows a similar argument to that of Lemma˜17. Details are omitted for brevity. ∎
D.2 Over-parameterized regression: implementation details
The implementation details are as follows. For IR-SCG, presented in Algorithm˜1, we set $\sigma_{t}=\varsigma(t+1)^{-1/4}$, where $\varsigma=10$, along with $\alpha_{t}=2/(t+2)$. For IR-FSCG, presented in Algorithm˜2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q\}+1)^{-1/2}$, where $\varsigma=10$, along with $\alpha_{t}=2/(t+1)$ for every $t\geq q$ and $\alpha_{t}=\log(q)/q$ for every $t<q$. Our linear minimization oracle involves optimization over an $\ell_{1}$-norm ball, which admits an analytical solution; see [26, Section 4.1]. We compare our proposed methods against SBCGI with stepsize $\gamma_{t}=0.01/(t+1)$, SBCGF with constant stepsize $\gamma_{t}=10^{-5}$ (both using $K_{t}=10^{-4}/\sqrt{t+1},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-7}/(t+1)^{3/4}$, $\rho_{t}=10^{3}(t+1)^{1/4}$, $r=0.5$), and SDBGD with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=10^{-6}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries or when the computation time exceeded 100 seconds, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{1}$-norm ball intersecting with a half-space is required, which we employ CVXPY [11] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [13, Algorithm 1]. To approximate the outer- and inner-level optimal values $F_{\operatorname{opt}},G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem and the reformulation (2) of the bilevel problem.
D.3 Dictionary learning: implementation details
In our experiment, we obtain $A,A^{\prime},\hat{X}$ from the code provided by Cao et al. [4] with $n=n^{\prime}=250$ and $p=40,q=50$ . In addition, we also choose $\delta=3$. For IR-SCG, presented in Algorithm˜1, we set $\sigma_{t}=\varsigma(t+1)^{-2/7}$ and $\alpha_{t}=(t+1)^{-6/7}$, where $\varsigma=0.1$. For IR-FSCG, presented in Algorithm˜2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-1/2}$, where $\varsigma=0.1$, along with $\alpha_{t}=(t+1)^{-3/4}$ for every $t\geq q+1$ and $\alpha_{t}=\log(q+1)/(q+1)$ for every $t\leq q$. In both experiments, the linear optimization oracle over the $\ell_{2}$-norm ball admits an analytical solution; see [26, Section 4.1]. For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.1(t+1)^{-2/3}$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-3}$ (both using $K_{t}=0.01(t+1)^{-1/3},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [27] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-4}/(t+1)^{3/4}$, $\rho_{t}=(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [21] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=5\times 10^{-3}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{2}$-norm ball intersecting with a half-space is required, which we employ CVXPY [11] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [13, Algorithm 1] to compute projection on $\ell_{1}$ norm ball and projection onto $\ell_{2}$ norm ball admits an analytical solution. To approximate inner-level optimal value $G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem.